### v2

@@ 2.4 Experimental validation
To evaluate learning performance, an 18-layer deep ResNet model was trained and tested on the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) dataset [26]. This dataset spans 1000 object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images [26].
@@
All models and algorithms were implemented in PyTorch [27] and run on a compute cluster using Nvidia A100 GPUs and Intel Xeon Platinum 8360Y processors. To initialize the weights of our models, we set $\mathbf{R}$ to the identity matrix and used He initialization [28] for $\mathbf{W}$. Models were trained for 50 epochs to minimize the categorical cross-entropy loss. Exploratory analysis revealed negligible performance difference as a function of batch size and a batch size of 256 was chosen. Instead of using the full batch of images in each decorrelation update step, we reduce the number of samples to 10%. We empirically found this value to show negligible loss in performance, while significantly speeding up runtime. The accuracy reported is the top-1 performance of the models. Reported wall-clock time was measured as the runtime of the training process, excluding performance testing.
Performance of DBP using either decorrelation ($\kappa=0$) or whitening ($\kappa=0.5$) was compared against regular backpropagation. For updating $\mathbf{W}$, we use the Adam optimizer [29] with a learning rate of $\eta=1.6\cdot 10^{-4}$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We added a small value of $10^{-8}$ to the denominator of the updates for numerical stability. The decorrelation matrix $\mathbf{R}$ is updated with stochastic gradient descent using a learning rate of $\epsilon=10^{-5}$. To ensure a fair comparison between algorithms, we made sure that all algorithms attained their highest convergence speed by performing a two-dimensional grid search over the BP ($\eta$) and decorrelation ($\epsilon$) learning rates. See Section B for the grid-search results.
@@ 3 Results
In the following, we analyse the performance of BP versus DBP on the 18-layer ResNet model trained on the ImageNet task.
@@ 3.1 DBP effectively decorrelates inputs to all network layers
In Fig. 3, the change is input correlation is shown for each layer during the training process for DBP. To quantify input correlation we use the mean squared values of the strictly lower triangular part of the $\mathbf{C}$ matrix. It can be seen that layer inputs become rapidly decorrelated during the first 10 epochs and correlations remain consistently low. This demonstrates that the decorrelating learning rule indeed effectively learns to decorrelate the inputs to all network layers.
Figure 3: Input decorrelation when training a ResNet-18 model on ImageNet for 50 epochs. Network layers are ordered from left to right and from top to bottom. Panel titles indicate the layer type.
@@ 3.2 DBP converges much faster than BP
We observe a significant improvement when training the ResNet model using DBP compared to regular BP. As shown in Fig. 4, DBP’s train accuracy (Fig. 4a) improves much faster than that of BP, indicating more effective training (computational work per epoch). The same observation holds for test accuracy (Fig. 4b). Additionally, DBP’s test accuracy (55.2%) also peaks above BP’s test accuracy (54.1%), while doing so in roughly half the number of epochs. Finally, we observe that a slightly higher test accuracy is achieved for whitening ($\kappa=0.5$) compared to decorrelation ($\kappa=0$), demonstrating the added benefit of whitening for this task.
The same patterns can be observed for the loss curves in Appendix C. Furthermore, as shown in Appendix D, a slightly lower test loss at equal train loss is obtained for DBP compared to BP, which shows that this does not come at the expense of generalization performance.
Figure 4: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train accuracy as a function of the number of epochs. b) Test accuracy as a function of the number of epochs.
3.3 DBP training can be achieved at shorter wall-clock times than BP
Recall that DBP requires additional computation in each epoch due to application of the decorrelation update rule in Eq. 10. This implies that faster convergence as a function of the number of epochs may not necessarily translate into more efficient training. Therefore, we also compare wall-clock times in Fig. 5.
Figure 5: Performance of BP and DBP using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train accuracy as a function of wall-clock time. b) Test accuracy as a function of wall-clock time.
[…]
3.4 DBP significantly reduces carbon emission
Figure 6 visualizes what the efficiency gains are when using DBP ($\kappa=0.5$) over BP in terms of carbon emission.
Figure 6: Efficiency gains when using DBP over BP. a) Training time required relative to BP’s time to maximal accuracy to reach a specific percentage of the maximum performance of DBP. b) Estimated reduction in carbon consumption when using DBP over BP. Estimate produced using https://github.com/mlco2/codecarbon. Dashed line indicates the moment when performance is equal to maximal BP test accuracy.
[…]
Figure 6b provides a measurement of the associated reduction in carbon emission if we run the model to achieve a specific percentage of the DBP test accuracy. The figure shows that a reduction of 640 grams of carbon emission is achieved if we train DBP until it reaches 100% of the BP test accuracy. If we run DBP until it reaches 100% of the DBP test accuracy, we still achieve a reduction of 418 grams of carbon emission.
@@ 4 Discussion
Increasing the efficiency of deep learning is of key importance if we are to reduce the carbon footprint of AI. In this paper, we have shown that decorrelated backpropagation provides a viable path towards more efficient deep learning. Results show that, by replacing BP with DBP, we can achieve higher test accuracy using about half the number of training epochs. Results also show that both decorrelation ($\kappa=0$) and whitening ($\kappa=0.5$) provide similar performance gains, with whitening yielding slightly better results. This does not come at the expense of generalization performance, as shown in Fig. 9. We also demonstrated a substantial reduction in carbon emission when using DBP over BP when training deep neural networks. Global carbon emissions can therefore be substantially reduced when applying DBP at scale.
These results indicate that performance is maximized for some optimal value of $\kappa$ which most effectively balances the covariance and variance constraints. This optimal value is task-dependent. Appendix F shows that training a medium-sized neural network on the CIFAR10 task yields similar performance gains but now the optimal performance is obtained for $\kappa=0$.
It should be mentioned that network-wide decorrelation comes with a number of additional advantages, which we did not further explore here. As shown in [17], decorrelated inputs allow for rapid linear approximation of non-linear functions, allowing for filter visualisation in deep networks with applications in explainable AI [30, 31]. It also allows for network compression at virtually no computational overhead, which may further decrease carbon consumption at inference time [13]. Intriguingly, input decorrelation and whitening have also been shown to be a feature of neural processing [32, 25, 33, 34, 35, 36, 37]. The work presented here suggests that decorrelation may be an important factor when considering synaptic plasticity mechanisms.
A number of considerations do need to be taken into account when using DBP. First, DBP requires updating and storing of decorrelation matrices needed for input decorrelation. Hence, the algorithm introduces some computational overhead and increased memory requirements. This is reflected by the increased wall-clock time per epoch. To mitigate this, more effective use could be made of the sparseness structure of the decorrelation matrix. For instance, rather than learning a full matrix $\mathbf{R}$ we may choose to learn a lower-triangular matrix, which can be achieved by using the strictly lower triangular part of $\mathbf{C}$ in Eq. 10. Other approaches to reduce wall-clock time may be the use of low-rank approximations of $\mathbf{R}$, using higher layer-specific decorrelation learning rates, or only incorporating decorrelation in that subset of the layers which have the biggest impact in terms of reducing convergence time [19].
Second, DBP needs proper fine-tuning since the stability of the algorithm depends on the two learning rates $\eta$ and $\epsilon$, whose ratio needs to be set appropriately such that the decorrelation rate throughout all network layers remains optimal. We empirically observe that this fine-tuning depends on the employed dataset, network architecture and loss function. Additional work on the interaction between these factors may provide insights into how to optimally set these parameters.
@@
An important consideration here is scaling the updates according to layer size. Note that Eq. 10 ignores the layer size in the normalisation, which implies that decorrelation updates have a larger impact in larger layers whereas whitening induced by $\kappa>0$ has a smaller impact in larger layers. In simulation work, we observe that such normalisation ensures that the impact of decorrelation updates is comparable for different layer sizes. However, in our empirical work using very deep neural networks, we find that ignoring this normalisation actually improves performance. There are different reasons why this may be the case. First, the resulting stronger decorrelation in deep large layers when ignoring normalisation may have a beneficial impact since deeper layers are more strongly affected by multiple preceding nonlinearities. Second, the decorrelation loss that is minimized by Eq. 10 may not be the optimal metric to ensure fastest convergence. That is, there may be a discrepancy between the decorrelation gradient direction and the optimal gradient direction to align the inputs.
@@
In this work, we demonstrated a significant speedup in deep residual networks, which are state-of-the-art models for computer vision. In our follow-up work, we aim to explore the efficiency gains that can be obtained for other computational tasks and network architectures. Exciting prospects are the use of decorrelated backpropagation for deep reinforcement learning [38, 39] or training of foundation models [40], both of which are notoriously resource intensive. We further expect that additional theoretical work on the optimal alignment of layer inputs combined with efficient low-level implementations will yield even larger gains in convergence speed. Concluding, by demonstrating the efficiency of decorrelated backpropagation in modern deep neural networks, we hope to contribute to reducing the energy consumption of modern AI systems.
Code availability
Code for training and evaluating the decorrelated backpropagation algorithm can be downloaded from https://github.com/artcogsys/decorbp.
@@ Appendix A Comparison with zero-phase component analysis
Zero-phase component analysis (ZCA) [24, 25] is defined by the whitening transform $\mathbf{x}=\mathbf{C}\mathbf{x}_{c}$, where $\mathbf{x}_{c}=\bar{\mathbf{x}}-\bm{\mu}$ with $\bm{\mu}$ the mean over datapoints and $\mathbf{C}$ is the whitening transform. Let $\mathbf{X}$ be the $n\times d$ matrix consisting of $n$ $d$-dimensional centered inputs. ZCA uses
@@
with $\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ the eigendecomposition of the covariance matrix $\mathbf{\Sigma}=\mathbf{X}\mathbf{X}^{\top}/n$.
Desjardins et al. [20] propose to periodically (during training) measure the correlation, $\mathbf{C}$, at every layer of a deep neural network and to then apply eigendecomposition method to compute the ZCA transform. This approach to achieving an exact whitening transform is an alternative to what we attempt in this work but has a number of drawbacks. First, the complexity of computing the eigendecomposition itself has a cost equivalent to a matrix multiplication - meaning that this step alone is equally expensive to our decorrelation method. Second, atop this decomposition, Desjardins et al. [20] further compute a matrix inverse of this transform which allows them to periodically compute an exact ZCA-based decorrelation within each layer of their network while also undoing its impact on the layer’s computation (by modifying the layer’s weight matrix $\mathbf{W}$) to ensure that the newly introduced transform does not change the network’s output. We avoid the second step by simultaneously optimising for both the decorrelation and task at every layer and at every update step, without directly computing the ZCA transform. This provides the benefit of not requiring any matrix inversion computation and also ensures that we are continuously optimising the decorrelation transform rather than doing so at fixed intervals.
Appendix B Hyper-parameter optimization
A grid search across learning rate hyper-parameters was performed to ensure that each algorithm reached its optimal performance. An exploratory analysis revealed that although $3.2\cdot 10^{-4}$ gave slightly better performance after five epochs for BP and DBP, later in training a learning rate of $1.6\cdot 10^{-4}$ was better. Therefore, the latter was chosen for all algorithms, which had the convenient side effect of using the same learning rate for all algorithms, making comparison more straightforward. Figure 7 depicts the grid search outcomes.

### v4

@@ 2.4 Experimental validation
To evaluate learning performance in DNNs, 18-, 34- and 50-layer ResNet models as well as an AlexNet model (Krizhevsky et al., 2012) were trained using both BP and DBP. Performance was evaluated using the ImageNet Large Scale Visual Recognition Challenge (ILSVRC) dataset (Russakovsky et al., 2015). This dataset spans 1000 object classes and contains 1,281,167 training images, 50,000 validation images and 100,000 test images.
@@
All models and algorithms were implemented in PyTorch (Paszke et al., 2019) and run on a compute cluster (Surf Snellius) using Nvidia A100 GPUs and Intel Xeon Platinum 8360Y processors. To initialize the weights of our models, we set $\mathbf{R}$ to the identity matrix and used He initialization (He et al., 2015) for $\mathbf{W}$. Models were trained to minimize the categorical cross-entropy loss. Exploratory analysis revealed negligible performance difference as a function of batch size and a batch size of 256 was chosen. Instead of using the full batch of images in each decorrelation update step, we reduce the number of samples to 10%. We empirically found this value to show negligible loss in performance, while significantly speeding up runtime. See Appendix B for an analysis. The accuracy reported is the top-1 performance of the models. Reported wall-clock time was measured as the runtime of the training process, excluding performance testing. Total compute time used is estimated at 400 hours, which includes initial exploration, hyper-parameter search and running the large-scale experiments.
For the $\kappa$ parameter, a value of 0.5 was chosen, balancing the covariance and variance constraints. For updating $\mathbf{W}$, we use the Adam optimizer (Kingma and Ba, 2014) with a learning rate of $\eta=1.6\cdot 10^{-4}$ and optimizer parameters $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We added a small value of $10^{-8}$ to the denominator of the updates for numerical stability. The decorrelation matrix $\mathbf{R}$ is updated with stochastic gradient descent using a learning rate of $\epsilon=10^{-5}$. To ensure a fair comparison between algorithms, we made sure that all algorithms attained their highest convergence speed by performing a two-dimensional grid search over the BP ($\eta$) and decorrelation ($\epsilon$) learning rates. See Appendix C for the grid search results on the ResNet18 model.
@@ 3 Results
In the following, we analyze the performance of DBP on different DNN models trained on the ImageNet task.
@@ 3.1 DBP effectively decorrelates inputs to all network layers
We first verified that DBP indeed effectively reduces input decorrelation across layers in DNNs. Figure 3 shows the change is input correlation for each layer of the ResNet18 model during the training process for DBP. To quantify input correlation we use the mean squared values of the strictly lower triangular part of the $\mathbf{C}$ matrix. Layer inputs become rapidly decorrelated during the first 10 epochs and correlations remain consistently low. This demonstrates that the decorrelating learning rule indeed effectively learns to decorrelate the inputs to all network layers.
Figure 3: Input decorrelation when training a ResNet18 model on ImageNet for 50 epochs. Network layers are ordered from left to right and from top to bottom. Panel titles indicate the layer type.
@@ 3.2 DBP converges much faster than BP
Figure 4 compares training and test convergence using DBP and BP as a function of the number of epochs across models. For each model, convergence of both train and test accuracy is much faster for DBP compared to BP, demonstrating the positive impact of decorrelation. Furthermore, test performance is consistently higher for DBP compared to BP.
Figure 4: Train and test performance of BP and DBP on Imagenet as a function of epochs for different deep neural network architectures. AlexNet was trained for 20 epochs and the deeper ResNet models were trained for 35 epochs. a) AlexNet. b) ResNet18. c) ResNet34. d) ResNet50.
3.3 DBP training yields shorter wall-clock times
As emphasized in the Introduction section, speedup in terms of epochs may not necessarily translate into speedup in terms of wall-clock time due to additional computational overhead. Figure 5 shows that, while the speedup is indeed attenuated due to this overhead, DBP remains faster than BP, reducing training time up to 50% across models when using peak test performance as a stopping criterion.
Figure 5: Train and test performance of BP and DBP on Imagenet as a function of wall-clock time for different deep neural network architectures. AlexNet was trained for 20 epochs and the deeper ResNet models were trained for 35 epochs. a) AlexNet. b) ResNet18. c) ResNet34. d) ResNet50.
3.4 DBP reduces carbon emission
Reduced wall-clock time does not only reduce the need for compute capacity but also decreases the carbon emission of deep learning. Figure 6 summarizes our results, showing faster learning at higher accuracy with reduced carbon consumption.
Figure 6: a) Maximum test accuracy b) Time needed to achieve maximum test accuracy c) Estimated reduction in carbon consumption when using DBP over BP. Estimates produced using https://github.com/mlco2/codecarbon.
3.5 Impact of whitening
The question remains how the choice of the whitening parameter $\kappa$ influences performance. To this end, we compared the performance of the ResNet18 model using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening) for three random initializations. Figure 7 show that performance is slightly improved for $\kappa=0.5$ compared to $\kappa=0$. However, as shown in Appendix D, the optimal choice of $\kappa$ depends on the chosen architecture and dataset.
Figure 7: Performance of BP and DBP on ResNet18 using $\kappa=0$ (decorrelation) and $\kappa=0.5$ (whitening). Reported results are the average of three randomly initialized networks, where minimal and maximal value are indicated by the error bars. a) Train accuracy as a function of the number of epochs. b) Test accuracy as a function of the number of epochs.
@@ 4 Discussion
Increasing the efficiency of deep learning is of key importance if we are to reduce the carbon footprint of AI. In this paper, using four deep neural network architectures trained on the Imagenet dataset, we have shown that decorrelated backpropagation, which decorrelates the inputs to each network layer, provides a viable path towards more efficient deep learning. Results show that, by replacing BP with DBP, we achieve higher test accuracy using up to half the number of training epochs, resulting in reduced carbon emission.
The observed reduction in training time and associated carbon emission of DBP is the result of a complex interplay between having a closer alignment of the gradient with the natural gradient, yielding faster convergence per epoch, and the need to decorrelate inputs across network layers, which comes with additional overhead per epoch. This overhead stems from a number of factors which apply to all $1\leq k\leq K$ network layers. First, the need to store the decorrelation matrices $\mathbf{R}_{k}$. Second, computing the gradient update (8) for each matrix $\mathbf{R}_{k}$ in parallel. Third, additional matrix multiplications $\mathbf{R}_{k}$ and $\mathbf{R}^{\top}_{k}$ during the forward and backward pass of the backpropagation algorithm. These factors contribute to increased memory and compute requirements though this is more than compensated for by the much faster convergence of DBP compared to BP.
These considerations suggest that even larger performance gains may be achievable if we can further reduce the memory and compute requirements of DBP. One way to achieve this may be to exploit sparseness structure in the decorrelation matrix. For instance, rather than learning a full matrix $\mathbf{R}$ we may choose to learn a lower-triangular matrix, which can be achieved by using the strictly lower triangular part of $\mathbf{C}$ in Eq. 8. Another approach may be to use a low-rank approximation of $\mathbf{R}$. Finally, we may choose to only incorporate decorrelation in that subset of layers which have the biggest impact in terms of reducing convergence time (Huangi et al., 2018). Further speedups may be achievable by improving the decorrelation algorithm itself, for instance, by using layer-specific decorrelation learning rates or subsampling rates that take layer properties into account. We expect that additional theoretical work on the optimal alignment of layer inputs combined with efficient low-level implementations will yield even greater gains in convergence speed.
It is important to realize that DBP needs proper fine-tuning. The stability of the algorithm depends on the interplay between the backpropagation learning rate $\eta$ and the decorrelation learning rate $\epsilon$. This ratio needs to be set appropriately such that the decorrelation rate throughout all network layers remains optimal. That is, a too slow decorrelation learning rate will not improve convergence speed while still incurring computational overhead. On the other hand, a too fast decorrelation learning rate will reduce the ability of backpropagation to keep up with changes in input statistics across the network, leading to instabilities. We empirically observe that this fine-tuning depends on the employed dataset, network architecture and loss function. Additional work on the interaction between these factors may provide insights into how to optimally set these parameters. Our results further show that performance is maximized for some optimal value of $\kappa$ which most effectively balances the covariance and variance constraints. As a rule of thumb, the setting $\kappa=0.5$ is reasonable though we may also choose to set $\kappa=0$, further simplifying the decorrelation learning rule to $\mathbf{R}\leftarrow\mathbf{R}-\epsilon\left\langle\mathbf{C}\right\rangle\mathbf{R}$.
@@
Another important consideration is normalizing the decorrelation update magnitude according to layer size. Note that Eq. 8 ignores the layer size in the normalization, which implies that decorrelation updates have a larger impact in larger layers whereas whitening induced by $\kappa>0$ has a smaller impact in larger layers. In simulation work, we observe that such normalization ensures that the impact of decorrelation updates is comparable for different layer sizes. However, in our empirical work using very deep neural networks, we find that disregarding this normalization actually improves performance. There are different reasons why this may be the case. First, the resulting stronger decorrelation in deep large layers when ignoring normalization may have a beneficial impact since deeper layers are more strongly affected by multiple preceding nonlinearities. Second, the decorrelation loss that is minimized by Eq. 8 may not be the optimal metric to ensure fastest convergence. That is, there may be an as yet unidentified discrepancy between the decorrelation gradient direction and the optimal gradient direction to align the inputs.
It should be mentioned that network-wide decorrelation comes with a number of additional advantages, which we did not further explore here. As shown in (Ahmad et al., 2023), decorrelated layer inputs allow for extremely efficient computation of a optimal linear mapping between arbitrary network layers. That is, when inputs are decorrelated, the matrix that needs to be inverted to compute that ordinary least squares solution becomes diagonal, reducing the inversion from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d)$ with $d$ the number of inputs. This property allows for filter visualization in deep networks with applications in explainable AI (Rudin, 2019; Ras et al., 2022) as well as for network compression at virtually no computational overhead, which may further decrease carbon consumption at inference time (Wang, 2021).
Intriguingly, input decorrelation and whitening have also been shown to be a feature of neural processing (Franke et al., 2017; Bell and Sejnowski, 1997; Pitkow and Meister, 2012; Segal et al., 2015; Dodds et al., 2019; Graham et al., 2006; King et al., 2013). The work presented here suggests that decorrelation may be an important factor when considering synaptic plasticity mechanisms, warranting further investigation.
@@
In this work, we demonstrated a significant speedup in deep feedforward neural networks, which are state-of-the-art models for computer vision. In follow-up work, we aim to explore the efficiency gains that can be obtained for other computational tasks and network architectures. Exciting prospects are the use of decorrelated backpropagation for deep reinforcement learning (Mnih et al., 2015; Schrittwieser et al., 2020) or training of generative models (Vaswani et al., 2017; Gu and Dao, 2024; Bommasani et al., 2021), both of which are notoriously resource intensive. Concluding, by demonstrating the efficiency of decorrelated backpropagation in modern deep neural networks, we hope to contribute to reducing the energy consumption of modern AI systems.
@@ Appendix A Comparison with zero-phase component analysis
Zero-phase component analysis (ZCA) [Bell and Sejnowski, 1995, 1997] is defined by the whitening transform $\mathbf{x}=\mathbf{C}\mathbf{x}_{c}$, where $\mathbf{x}_{c}=\bar{\mathbf{x}}-\bm{\mu}$ with $\bm{\mu}$ the mean over datapoints and $\mathbf{C}$ is the whitening transform. Let $\mathbf{X}$ be the $n\times d$ matrix consisting of $n$ $d$-dimensional centered inputs. ZCA uses
@@
with $\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$ the eigendecomposition of the covariance matrix $\mathbf{\Sigma}=\mathbf{X}\mathbf{X}^{\top}/n$. Desjardins et al. [2015] propose to periodically (during training) measure the correlation, $\mathbf{C}$, at every layer of a deep neural network and to then apply eigendecomposition method to compute the ZCA transform.
This approach to achieving an exact whitening transform is an alternative to what we attempt in this work but has a number of drawbacks. First, the complexity of computing the eigendecomposition itself has a cost equivalent to a matrix multiplication - meaning that this step alone is equally expensive to our decorrelation method. Second, atop this decomposition, Desjardins et al. [2015] further compute a matrix inverse of this transform which allows them to periodically compute an exact ZCA-based decorrelation within each layer of their network while also undoing its impact on the layer’s computation (by modifying the layer’s weight matrix $\mathbf{W}$) to ensure that the newly introduced transform does not change the network’s output. We avoid the second step by simultaneously optimising for both the decorrelation and task at every layer and at every update step, without directly computing the ZCA transform. This provides the benefit of not requiring any matrix inversion computation and also ensures that we are continuously optimising the decorrelation transform rather than doing so at fixed intervals.
Appendix B Impact of subsampling
Figure 8 shows test performance as a function of wall-clock time for AlexNet trained on ImageNet at different sampling frequencies and sampling rates. Sampling 10% of the batch appears sufficient for robust decorrelation while keeping computational overhead limited. Trying to save even more computation time by only updating the decorrelation matrix every 5 or 10 batches does not appear to speed up convergence in this experiment. Based on these results, a sampling rate of 10% was used in all experiments, while the sampling frequency was not adjusted.
Figure 8: AlexNet test accuracy on ImageNet a) Sampling every batch, 5 batches, and 10 batches. b) Sampling 100%, 50%, and 10% of the batch.
Appendix C Hyper-parameter optimization
A grid search across learning rate hyper-parameters was performed to ensure that each algorithm reached its optimal performance. An exploratory analysis revealed that although $3.2\cdot 10^{-4}$ gave slightly better performance after five epochs for BP and DBP, later in training a learning rate of $1.6\cdot 10^{-4}$ was better. Therefore, the latter was chosen for all algorithms, which had the convenient side effect of using the same learning rate for all algorithms, making comparison more straightforward. Figure 9 depicts the grid search results for the ResNet18 model.